The extension in a string (obeying Hooke's law)  is x. The speed of sound in the stretched string is v. If the extension in the string is increased to 1.5x, what would the speed of sound be?

Grade:12

4 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

v/x=v'/1.5x

v'=1.5v

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi



sapna .
11 years ago

acc to me the correct answer is 1.22v as v is proportional to √x.

Maurya Sharma
5 years ago
Y=Fl/A∆l
 
F=(YA/l) ∆l 
F js Directly proportional to change in length
 
According to wave motion, v is Directly proportional to √F
 
Therefore, v is directly proportional to √∆l
 
V=kx 
V'=k√1.5x
V'=1.22v
 
 
 
 
 
ankit singh
3 years ago
 
 
Y=Fl/A∆l
 
F=(YA/l) ∆l 
F js Directly proportional to change in length
 
According to wave motion, v is Directly proportional to √F
 
Therefore, v is directly proportional to √∆l
 
V=kx 
V'=k√1.5x
V'=1.22v